# Semilinear spaces of matrices with transitivity, rank or spectral constraints

### v1

@@ 1.4 Main new results on trivial spectrum subspaces
The integers $n_{1},\dots,n_{p}$ are uniquely determined by $n$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ the matrix $P_{i}$ is uniquely determined by $\mathcal{M}$ up to star-congruence and multiplication by an element of $\mathbb{F}^{\times}$.
Conversely, for any partition $n=n_{1}+\cdots+n_{p}$ and any list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{D})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$ of $e$-nonisotropic matrices, the space $P_{1}^{-1}\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee P_{p}^{-1}\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D})$, as well as all its conjugates in $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, is an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
@@ 1.5 The strategy, and intransitive operator spaces
At this point, we have stated the main results of the present article, which are the ones on trivial spectrum spaces. They are far from the only results here, and we will expand on the other results in a moment.
But before we come to the other results, let us explain the main strategy for proving those on trivial spectrum spaces. So far, we have not succeeded in generalizing the techniques of [15, 18] to the case $d>1$. This has led us to reconsider an alternative proof of Theorem 1.4, limited to fields with more than n1n-1 elements, which we discovered ten years ago [19]. Critical to the proof is the observation that the $\mathbb{F}$-trivial spectrum property implies a lack of transitivity from the said spaces of operators, meaning that if $\mathcal{S}$ has $\mathbb{F}$-trivial spectrum then $x\not\in\mathcal{S}x$ for every nonzero vector $x\in V\smallsetminus\{0\}$, and in particular $\mathcal{S}x\subsetneq V$. The key insight is then to introduce the dual operator space $\widehat{\mathcal{S}}$ of $\mathcal{S}$, which consists of all the evaluation mappings
$\widehat{x}:u\in\mathcal{S}\mapsto u(x)\in V,\quad\text{with $x\in V$.}$
Then $\widehat{\mathcal{S}}$ is a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(\mathcal{S},V)$, and every element of $\widehat{\mathcal{S}}$ is nonsurjective and hence has its rank less than $\dim_{\mathbb{F}}V$. The key in [19] was to rely upon the theory of primitive spaces of bounded rank matrices, studied first by Atkinson and Lloyd [1, 2] and rediscovered a few years later by Eisenbud and Harris [4]. Thanks to Atkinson’s classification of primitive matrix spaces that satisfy an extremal condition between the numbers of rows and columns (theorem C of [2]), and more precisely to our own generalization to semi-primitive spaces (see section 5 of [21] and a restatement as theorem 3.1 in [19]), it was then possible to find an easy derivation of Theorem 1.4 in case $|\mathbb{F}|\geq n$ (an unavoidable assumption if one wants to use Atkinson’s theorem).
Yet, for the case $d>1$ a major issue is that the space $\widehat{\mathcal{S}}$ is far from satisfying Atkinson’s extremal condition between the respective dimensions of the source and target spaces, even if $\mathcal{S}$ is optimal; hence nothing can be obtained from the theory of primitive spaces of bounded rank matrices. Nevertheless, a version of Atkinson’s theorem (theorem 5.3 of [21]) can be entirely reformulated in the viewpoint of spaces of operators that lack transitivity, as we shall now see.
@@
Notation 1.4.
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $V^{\prime}$ be a $\mathbb{D}$-linear subspace of $V$. Given a subset $\mathcal{S}$ of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, we denote by $\mathcal{S}^{V^{\prime}}:=\mathcal{S}\cap\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V^{\prime})$ the subset consisting of the elements of $\mathcal{S}$ that map into $V^{\prime}$.
@@
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. We say that $\mathcal{S}$ is:
  • intransitive whenever $\forall x\in U,\;\mathcal{S}x\neq V$ ;
  • deeply intransitive whenever it is intransitive and, for every nonzero linear subspace $V^{\prime}$ of $V$, the set $\mathcal{S}^{V^{\prime}}$ is intransitive as an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V^{\prime})$.
In any case, the transitive rank $\operatorname{\operatorname{trk}}(\mathcal{S})$ of $\mathcal{S}$ is defined as the greatest possible dimension $\dim_{\mathbb{F}}(\mathcal{S}x)$ when $x$ ranges over $U$.
Note that $\mathcal{S}$ is intransitive only if $V\neq\{0\}$, and in the definition of deep intransitivity we can replace the condition that $\mathcal{S}$ be intransitive with the one that $V\neq\{0\}$.
It is obvious that if $\mathcal{S}$ is deeply intransitive then so is $\mathcal{S}^{V^{\prime}}$ for every nonzero $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$. Constructing large deeply intransitive spaces can be achieved by using nonzero bilinear forms, as we shall now explain.
@@
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. An $\mathbb{F}$-bilinear form $b:U\times V\rightarrow\mathbb{F}$ is called an alternator of $\mathcal{S}$ whenever
$\forall x\in U,\;\forall f\in\mathcal{S},\;b(x,f(x))=0.$
The set of all such forms is denoted by $\operatorname{\operatorname{Alt}}(\mathcal{S})$ (it is obviously a linear subspace of the set of all $\mathbb{F}$-bilinear forms on $U\times V$).
@@
Let $U$ and $V$ be right-$\mathbb{D}$-vector spaces and $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
@@
    If $\mathcal{S}$ has a nonzero alternator then $\mathcal{S}$ is intransitive.
@@
    If $\mathcal{S}$ has a right-nondegenerate alternator and $V\neq\{0\}$ then $\mathcal{S}$ is deeply intransitive.
@@
Assume that $\mathcal{S}$ has a nonzero alternator $b$. It follows by polarizing that $\forall(x,x^{\prime})\in U^{2},\;\forall f\in\mathcal{S},\;b(x,f(x^{\prime}))+b(x^{\prime},f(x))=0$, from which we observe that all the elements of $\mathcal{S}$ map $\operatorname{\operatorname{Lrad}}(b)$ into $\operatorname{\operatorname{Rrad}}(b)$. The latter is not the full space $V$ because $b$ is nonzero. Now, let $x\in U$. If $x\in\operatorname{\operatorname{Lrad}}(b)$ then $\mathcal{S}x\subseteq\operatorname{\operatorname{Rrad}}(b)\subset V$. Otherwise $\mathcal{S}x$ is included in $\operatorname{\operatorname{Ker}}b(x,-)$, an $\mathbb{F}$-linear hyperplane of $V$. Hence $\mathcal{S}$ is intransitive.
Assume now that $\mathcal{S}$ has a right-nondegenerate alternator $b$ and $V\neq\{0\}$. Let $V^{\prime}$ be a nonzero $\mathbb{D}$-linear subspace of $V$. The restriction $b^{\prime}$ of $b$ to $U\times V^{\prime}$ is nonzero because $b$ is right-nondegenerate, and obviously $b^{\prime}$ is an alternator of $\mathcal{S}^{V^{\prime}}$. Hence $\mathcal{S}^{V^{\prime}}$ is intransitive, by point (a). ∎
Hence, whenever we take a right-nondegenerate bilinear form $b:U\times V\rightarrow\mathbb{F}$ and $V\neq\{0\}$, every $\mathbb{F}$-linear subspace of $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive. Atkinson’s theorem can be essentially summed up in saying that all large enough operator spaces are of this kind. Let us now be more precise:
@@
Let $U$ and $V$ be finite-dimensional $\mathbb{F}$-vector spaces and $\mathcal{S}$ be a deeply intransitive linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$. Set $n:=\dim V$ and assume that $|\mathbb{F}|\geq n$. Then:
@@
    $\dim\mathcal{S}\leq\frac{n(n-1)}{2}$, and even $\dim\mathcal{S}\leq\frac{(n-1)(n-2)}{2}$ if $\operatorname{\operatorname{trk}}(\mathcal{S})<n-1$.
@@
    If $\dim\mathcal{S}\geq\frac{n(n-1)}{2}-(n-3)$ then $\mathcal{S}$ has a right-nondegenerate alternator $b$, and its alternator space $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$.
As an extension of point (b), we note, with the same polarization argument as in the proof of Proposition 1.18, that all the operators in $\mathcal{S}$ then vanish on the left radical $\operatorname{\operatorname{Lrad}}(b)$, and hence $\mathcal{S}$ can naturally be identified with a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U/\operatorname{\operatorname{Lrad}}(b),V)$, for which the form induced by $b$ on $(U/\operatorname{\operatorname{Lrad}}(b))\times V$ is a nondegenerate alternator. This fundamentally reduces the situation to the one where $\dim U=\dim V$, in which case we could go further and simply consider spaces of endomorphisms of $V$.
We will now make several comments on Atkinson’s theorem. First of all, we must state that Atkinson’s formulation of the theorem has several additional assumptions that turn out to be useless. Atkinson frames his result in the language of matrix spaces with bounded rank, and makes an assumption of primitivity. Primitivity consists of as many as four conditions that roughly mean that the rank constraint cannot be easily localized. In contrast, here we replace all four assumptions in the definition of primitivity by a single assumption, which is deep intransitivity (plain intransitivity is of course a far too weak hypothesis, allowing, e.g., $\mathcal{S}$ to be the space of all operators from $U$ to a given linear hyperplane of $V$). The deep intransitivity assumption is a dual counterpart of what Atkinson and Lloyd [1] called the row condition for matrix spaces with bounded rank. The row condition for such a matrix space $\mathcal{M}\subseteq\operatorname{\operatorname{M}}_{n,p}(\mathbb{F})$ means that, whenever we have invertible matrices $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$ and $Q\in\operatorname{\operatorname{GL}}_{p}(\mathbb{F})$, and integers $r\in\mathopen{[\![}1,n\mathclose{]\!]}$ and $s\in\mathopen{[\![}1,p\mathclose{]\!]}$ such that every matrix $M$ of $\mathcal{M}$ looks like
@@
Now, if we take an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$, consider its dual operator space $\widehat{\mathcal{S}}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(\mathcal{S},V)$, and represent the latter by a matrix space $\mathcal{M}$ in bases of $\mathcal{S}$ and $V$, the row condition on $\mathcal{M}$ is equivalent to the fact that whenever we take a nonzero linear subspace $\mathcal{S}^{\prime}$ of $\mathcal{S}$ and we know that all the operators in $\mathcal{S}^{\prime}$ map into some nonzero $\mathbb{F}$-linear subspace $V^{\prime}$ of $V$, then the rank of $s^{\prime}\in\mathcal{S}^{\prime}\mapsto s^{\prime}(x)\in V^{\prime}$ is less than the dimension of $V^{\prime}$ for all $x\in U$, which is exactly saying that $\mathcal{S}^{\prime}$ is intransitive as a linear subspace of mappings from $U$ to $V^{\prime}$. Hence, $\mathcal{M}$ satisfies the row condition if and only if $\mathcal{S}$ is deeply intransitive.
Among the four conditions that define primitivity, two are worth mentioning:
@@
An operator space $\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$ between $\mathbb{F}$-vector spaces $U$ and $V$ is called:
  • source-reduced when $\underset{f\in\mathcal{S}}{\bigcap}\operatorname{\operatorname{Ker}}f=\{0\}$, i.e., no nonzero vector is annihilated by all the operators in $\mathcal{S}$;
  • target-reduced when $\underset{f\in\mathcal{S}}{\sum}\operatorname{\operatorname{Im}}f=V$, i.e., no proper linear subspace of $V$ includes the range of all the operators in $\mathcal{S}$.
Finally, $\mathcal{S}$ is called reduced when it is both source-reduced and target-reduced.
Atkinson’s theorem is frequently formulated by adding the provision that the operator space $\mathcal{S}$ under consideration is reduced. However, this assumption is not necessary for the formulation in terms of intransitive spaces. Indeed, on the one hand the dual operator space $\widehat{\mathcal{S}}$ of an operator space $\mathcal{S}\subseteq\operatorname{\operatorname{Hom}}(U,V)$ is automatically source-reduced (as only the zero operator can be annihilated by all the evaluation mappings). On the other hand, if $\mathcal{S}$ is not target-reduced we can take the sum $V^{\prime}$ of all the ranges of the operators in $\mathcal{S}$, then observe that $\widehat{\mathcal{S}}$ can be viewed as a reduced subspace of $\operatorname{\operatorname{Hom}}(\mathcal{S},V^{\prime})$; applying point (a) in Atkinson’s theorem then readily shows that $\dim\mathcal{S}\leq\frac{\dim V^{\prime}(\dim V^{\prime}-1)}{2}\cdot$
Atkinson’s theorem deals with vector spaces over $\mathbb{F}$, but here we need to extend the situation to $\mathbb{F}$-linear subspaces of $\mathbb{D}$-linear mappings. In that context, it makes no sense to try to extend Atkinson’s formulation, as by representing the operators of $\widehat{\mathcal{S}}$ by matrices with entries in $\mathbb{F}$, we forget that the operators in $\mathcal{S}$ are $\mathbb{D}$-linear, and not only $\mathbb{F}$-linear. The correct way to generalize the results seems to be as a theorem on intransitive operator spaces. We now state the generalization:
@@
Let $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces and $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. Set $n:=\dim_{\mathbb{D}}V$ and assume that $|\mathbb{F}|\geq nd$. Then:
@@
    $\dim\mathcal{S}\leq\alpha(n,d)-n+\max(2-d,0)$ if $\operatorname{\operatorname{trk}}(\mathcal{S})<nd-1$.
@@
    If $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$ then $\mathcal{S}$ has a right-nondegenerate alternator $b$, and its alternator space is $1$-dimensional.
@@
Due to point (a), the inequality $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$ of point (c) requires that $n\geq 2$.
@@
Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. Set $n:=\dim_{\mathbb{D}}V$, assume that $|\mathbb{F}|\geq nd$, that $\mathcal{S}$ has a right-nondegenerate alternator $b$ and that $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$.
Then $\mathbb{D}$ has quadratic type over $\mathbb{F}$, and for any associated pair $(\sigma,e)$ the form $b$ equals $(x,y)\mapsto e(B(x,y))$ for a unique right-nondegenerate $\sigma$-sesquilinear form $B$ on $U\times V$, and of course $\mathcal{S}\subseteq\mathcal{A}_{b,\mathbb{D}}$.
@@
Having these results, it will be fairly easy to derive the ones on optimal irreducible trivial spectrum subspaces, by adapting the method of [19]. The main difficulty is that although a trivial spectrum subspace $\mathcal{S}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (with $V$ nonzero) is obviously intransitive (since $\mathcal{S}x$ can never contain $x$ when $x$ is a nonzero vector of $V$), it is however not deeply intransitive in general.

### v2

@@ 1.4 The new results on trivial spectrum subspaces
The integers $n_{1},\dots,n_{p}$ are uniquely determined by $n$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ the matrix $P_{i}$ is uniquely determined by $\mathcal{M}$ up to homo-star-congruence.
Conversely, for any partition $n=n_{1}+\cdots+n_{p}$ and any list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{D})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$ of $e$-nonisotropic matrices, the space $P_{1}^{-1}\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee P_{p}^{-1}\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D})$, as well as all its conjugates in $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, is an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$.
The corresponding results for the initial problem of optimal unital affine subspaces will be stated and proved in Section 7.
@@ 1.5 The strategy, and intransitive operator spaces
Let us now explain the main strategy for proving our new results on trivial spectrum spaces.
So far, we have not succeeded in generalizing the techniques of [10, 12] to the case $d>1$. This has led us to reconsider an alternative proof of Theorem 1.4, limited to fields with more than n1n-1 elements, which we discovered ten years ago [13]. Critical to the proof is the observation that the trivial spectrum property implies a lack of transitivity from the said spaces of operators, meaning that if $S$ has trivial spectrum then $x\not\in Sx$ for every nonzero vector $x\in V\smallsetminus\{0\}$, and in particular $Sx\subsetneq V$. The key insight is then to introduce the dual operator space $\widehat{S}$ of $S$, which consists of all the evaluation mappings
$\widehat{x}:u\in S\mapsto u(x)\in V,\quad\text{with $x\in V$.}$
Then $\widehat{S}$ is a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(S,V)$, and every element of $\widehat{S}$ is nonsurjective and hence has its rank less than $\dim_{\mathbb{F}}V$. The key in [13] was to rely upon the theory of primitive spaces of bounded rank matrices, studied first by Atkinson and Lloyd [1, 2] and rediscovered a few years later by Eisenbud and Harris [3] (see also the recent [6]). Thanks to Atkinson’s classification of primitive matrix spaces that satisfy an extremal condition between the numbers of rows and columns (theorem C of [2]), and more precisely to our own generalization to semi-primitive spaces (see section 5 of [15] and a restatement as theorem 3.1 in [13]), it was then possible to find an easy derivation of Theorem 1.4 in case $|\mathbb{F}|\geq n$ (an unavoidable assumption if one wants to use Atkinson’s theorem).
Yet, for the case $d>1$ a major issue is that the space $\widehat{S}$ is far from satisfying Atkinson’s extremal condition between the respective dimensions of the source and target spaces, even if $S$ is optimal; hence nothing can be obtained from the theory of primitive spaces of bounded rank matrices. Nevertheless, we discovered that a version of Atkinson’s theorem (theorem 5.3 of [15]) can be entirely reformulated in the viewpoint of spaces of operators that lack transitivity, as we shall now see.
@@
Notation 1.5.
Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $V^{\prime}$ be a $\mathbb{D}$-linear subspace of $V$. Given a subset $S$ of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$, we denote by $S^{V^{\prime}}:=S\cap\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V^{\prime})$ the subset consisting of the elements of $S$ that map into $V^{\prime}$.
@@
Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. We say that $S$ is:
  • intransitive whenever $\forall x\in U,\;Sx\neq V$ ;
  • deeply intransitive whenever it is intransitive and, for every nonzero linear subspace $V^{\prime}$ of $V$, the set $S^{V^{\prime}}$ is intransitive as an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V^{\prime})$.
In any case, the transitive rank $\operatorname{\operatorname{trk}}(S)$ of $S$ is defined as the greatest possible dimension $\dim_{\mathbb{F}}(Sx)$ when $x$ ranges over $U$.
Note that $S$ is intransitive only if $V\neq\{0\}$, and in the definition of deep intransitivity we can replace the condition that $S$ be intransitive with the one that $V\neq\{0\}$.
It is obvious that if $S$ is deeply intransitive then so is $S^{V^{\prime}}$ for every nonzero $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$. Constructing large deeply intransitive spaces can be achieved by using nonzero bilinear forms, as we shall now explain.
@@
Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. An $\mathbb{F}$-bilinear form $b:U\times V\rightarrow\mathbb{F}$ is called an alternator of $S$ whenever
$\forall x\in U,\;\forall f\in S,\;b(x,f(x))=0.$
The set of all such forms is denoted by $\operatorname{\operatorname{Alt}}(S)$ (it is obviously a linear subspace of the set of all $\mathbb{F}$-bilinear forms on $U\times V$).
@@
Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$.
@@
    If $S$ has a nonzero alternator then $S$ is intransitive.
@@
    If $S$ has a right-nondegenerate alternator and $V\neq\{0\}$ then $S$ is deeply intransitive.
@@
Assume that $S$ has a nonzero alternator $b$. It follows by polarizing that $\forall(x,x^{\prime})\in U^{2},\;\forall f\in S,\;b(x,f(x^{\prime}))+b(x^{\prime},f(x))=0$, from which we observe that all the elements of $S$ map $\operatorname{\operatorname{Lrad}}(b)$ into $\operatorname{\operatorname{Rrad}}(b)$. The latter is not the full space $V$ because $b$ is nonzero. Now, let $x\in U$. If $x\in\operatorname{\operatorname{Lrad}}(b)$ then $Sx\subseteq\operatorname{\operatorname{Rrad}}(b)\subset V$. Otherwise $Sx$ is included in $\operatorname{\operatorname{Ker}}b(x,-)$, an $\mathbb{F}$-linear hyperplane of $V$. Hence $S$ is intransitive.
Assume now that $S$ has a right-nondegenerate alternator $b$ and $V\neq\{0\}$. Let $V^{\prime}$ be a nonzero $\mathbb{D}$-linear subspace of $V$. The restriction $b^{\prime}$ of $b$ to $U\times V^{\prime}$ is nonzero because $b$ is right-nondegenerate, and obviously $b^{\prime}$ is an alternator of $S^{V^{\prime}}$. Hence $S^{V^{\prime}}$ is intransitive, by point (a). ∎
Hence, whenever we take a right-nondegenerate bilinear form $b:U\times V\rightarrow\mathbb{F}$ and $V\neq\{0\}$, every $\mathbb{F}$-linear subspace of $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive. When $\mathbb{D}=\mathbb{F}$, Atkinson’s theorem can be essentially summed up in saying that all large enough operator spaces are of this kind. Let us now be more precise:
@@
Let $U$ and $V$ be finite-dimensional $\mathbb{F}$-vector spaces and $S$ be a deeply intransitive linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$. Set $n:=\dim V$ and assume that $|\mathbb{F}|\geq n$. Then:
@@
    $\dim S\leq\frac{n(n-1)}{2}$, and even $\dim S\leq\frac{(n-1)(n-2)}{2}$ if $\operatorname{\operatorname{trk}}(S)<n-1$.
@@
    If $\dim S\geq\frac{n(n-1)}{2}-(n-3)$ then $S$ has a right-nondegenerate alternator $b$, and $\operatorname{\operatorname{Alt}}(S)=\mathbb{F}b$.
As an extension of point (b), we note, with the same polarization argument as in the proof of Proposition 1.20, that all the operators in $S$ then vanish on the left radical $\operatorname{\operatorname{Lrad}}(b)$, and hence $S$ can naturally be identified with a linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U/\operatorname{\operatorname{Lrad}}(b),V)$, for which the form induced by $b$ on $(U/\operatorname{\operatorname{Lrad}}(b))\times V$ is a nondegenerate alternator. This fundamentally reduces the situation to the one where $\dim U=\dim V$, in which case we could go further and simply consider spaces of endomorphisms of $V$.
We will now make several comments on Atkinson’s theorem. First of all, we must state that Atkinson’s formulation of the theorem has several additional assumptions that turn out to be useless. Atkinson frames his result in the language of matrix spaces with bounded rank, and makes an assumption of primitivity. Primitivity consists of as many as four conditions that roughly mean that the rank constraint cannot be easily localized. In contrast, here we replace all four assumptions in the definition of primitivity by a single assumption, which is deep intransitivity (plain intransitivity is of course a far too weak hypothesis, allowing, e.g., $S$ to be the space of all operators from $U$ to a given linear hyperplane of $V$). The deep intransitivity assumption is a dual counterpart of what Atkinson and Lloyd [1] called the row property for matrix spaces with bounded rank. The row property for such a matrix space $\mathcal{M}\subseteq\operatorname{\operatorname{M}}_{n,p}(\mathbb{F})$ means that, whenever we have invertible matrices $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$ and $Q\in\operatorname{\operatorname{GL}}_{p}(\mathbb{F})$, and integers $r\in\mathopen{[\![}1,n\mathclose{]\!]}$ and $s\in\mathopen{[\![}1,p\mathclose{]\!]}$ such that every matrix $M$ of $\mathcal{M}$ looks like
@@
Now, if we take an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$, consider its dual operator space $\widehat{S}\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(S,V)$, and represent the latter by a matrix space $\mathcal{M}$ in bases of $S$ and $V$, the row condition on $\mathcal{M}$ is equivalent to the fact that whenever we take a nonzero linear subspace $S^{\prime}$ of $S$ and we know that all the operators in $S^{\prime}$ map into some nonzero $\mathbb{F}$-linear subspace $V^{\prime}$ of $V$, then the rank of $s^{\prime}\in S^{\prime}\mapsto s^{\prime}(x)\in V^{\prime}$ is less than the dimension of $V^{\prime}$ for all $x\in U$, which is exactly saying that $S^{\prime}$ is intransitive as a linear subspace of mappings from $U$ to $V^{\prime}$. Hence, $\mathcal{M}$ has the row property if and only if $S$ is deeply intransitive.
Among the four conditions that define primitivity under the Atkinson-Lloyd definition, two are worth mentioning:
@@
An operator space $S\subseteq\operatorname{\operatorname{Hom}}_{\mathbb{F}}(U,V)$ between $\mathbb{F}$-vector spaces $U$ and $V$ is called:
  • source-reduced when $\underset{f\in S}{\bigcap}\operatorname{\operatorname{Ker}}f=\{0\}$, i.e., no nonzero vector is annihilated by all the operators in $S$;
  • target-reduced when $\underset{f\in S}{\sum}\operatorname{\operatorname{Im}}f=V$, i.e., no proper linear subspace of $V$ includes the range of all the operators in $S$.
Finally, $S$ is called reduced when it is both source-reduced and target-reduced.
Atkinson’s theorem is frequently formulated by adding the provision that the operator space $S$ under consideration is reduced (see e.g., [15]). However, this assumption is not necessary for the formulation in terms of intransitive spaces. Indeed, on the one hand the dual operator space $\widehat{S}$ of an operator space $S\subseteq\operatorname{\operatorname{Hom}}(U,V)$ is automatically source-reduced (as only the zero operator can be annihilated by all the evaluation mappings). On the other hand, if $S$ is not target-reduced we can take the sum $V^{\prime}$ of all the ranges of the operators in $S$, then observe that $\widehat{S}$ can be viewed as a reduced subspace of $\operatorname{\operatorname{Hom}}(S,V^{\prime})$; applying point (a) in Atkinson’s theorem then readily shows that $\dim S\leq\frac{\dim V^{\prime}(\dim V^{\prime}-1)}{2}\cdot$
Atkinson’s theorem deals with vector spaces over $\mathbb{F}$, but here we need to extend the situation to $\mathbb{F}$-linear subspaces of $\mathbb{D}$-linear mappings. In that context, it makes no sense to try to extend Atkinson’s formulation, as by representing the operators of $\widehat{S}$ by matrices with entries in $\mathbb{F}$, we miss the critical fact that the operators in $S$ are $\mathbb{D}$-linear, and not only $\mathbb{F}$-linear. The correct way to generalize the result seems to be as a theorem on intransitive operator spaces. We now state the generalization:
@@
Let $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces and $S$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. Set $n:=\dim_{\mathbb{D}}V$ and assume that $|\mathbb{F}|\geq nd$. Then:
@@
    $\dim S\leq\alpha(n,d)-n+\max(2-d,0)$ if $\operatorname{\operatorname{trk}}(S)<nd-1$.
@@
    If $\dim S\geq\alpha(n,d)-n+\max(4-d,2)$ then $S$ has a right-nondegenerate alternator $b$, and $\operatorname{\operatorname{Alt}}(S)=\mathbb{F}b$.
@@
Due to point (a), the inequality $\dim S\geq\alpha(n,d)-n+\max(4-d,2)$ of point (c) requires that $n\geq 2$, and even that $n\geq 3$ if $d=1$.
@@
Let $U$ and $V$ be $\mathbb{D}$-vector spaces and $S$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$. Set $n:=\dim_{\mathbb{D}}V$, assume that $|\mathbb{F}|\geq nd$, that $S$ has a right-nondegenerate alternator $b$ and that $\dim S\geq\alpha(n,d)-n+\max(4-d,2)$.
Then $\mathbb{D}$ has quadratic type over $\mathbb{F}$, and for any associated pair $(\sigma,e)$ the form $b$ equals $(x,y)\mapsto e(B(x,y))$ for a unique right-nondegenerate $\sigma$-sesquilinear form $B$ on $U\times V$, and of course $S\subseteq\mathcal{A}_{b,\mathbb{D}}$.
@@
Having these results, it will be fairly easy to derive the ones on optimal irreducible trivial spectrum subspaces, by adapting the method of [13]. The main difficulty is that although a trivial spectrum subspace $S$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (with $V$ nonzero) is obviously intransitive (since $Sx$ can never contain $x$ when $x$ is a nonzero vector of $V$), it is however not deeply intransitive in general.